Question 2

The manager of War Games has two potential games to develop, Nimrod and F1-11. The design teams have provided the following free cash flow forecasts. The cost of capital is 7%.  

Time period

Nimrod

F1-11

Initial Investment

£1,000,000

£1,200,000

Year 1

£300,000

£200,000

Year 2

£950,000

£600,000

Year 3

£650,000

£1,200,000

Year 4

£600,000

£1,400,000

  1. Calculate the Net Present Value of each project and advise the manager which project(s), he should accept.
  1. marks)

  1. Assuming that the manager has only £1,200,000 to invest.

(i)     Which project should he choose using the NPV method?

(ii)    Would the choice be different using the Payback method?

         

(10 marks)

      

  1. Explain why some managers prefer the payback method to NPV?

(5 marks)
Answer #1

1. Calculation of NPV of Nimrod

Year Cashflow Discount factor 7% Present cash flow
0 1,000,000 =1/(1+r)^n

(1,000,000)

1 300,000 1/(1+0.07)^1= 0.9345 280,350
2 950,000 1/(1+0.07)^2 = 0.8734 829,730
3 650,000 1/(1+0.07)^3= 0.8162 530,530
4 600,000 1/(1+0.07)^4= 0.7628 457,680
2,098,290

NPV = Present cash inflow - cash outflow

Net present cash Inflow = 2,098,290

Cash outflow = 1,000,000

NPV = 2,098,290 - 1,000,000 = 1,098,290

NPV calculation of F1-11

Year Cash flow Discounted value Present value cash flow
0 1,200,000 1/(1+0.07)^n (1,200,000)
1 200,000 =1/(1+0.07)^1=0.9345 186,900
2 600,000 =1/(1+0.07)^2=0.8734 524,040
3 1,200,000 =1/(1+0.07)^3=0.8162 979,440
4 1,400,000 =1/(1+0.07)^4=0.7628 1,067,920
2,758,300

NPV = Present value cash flow - cash outflow

= 2,758,300- 1,200,000 = 1,558,300

Manger should accept project F1-11 because it has higher NPV than Nimrod

b)

(i) Manager would choose NPV of higher value and that is F1-11

(ii) Calculation of Pay back Period of Nimrod

Year Cash flow Cumulative cashflow
1 300,000 300,000
2 950,000 1,250,000
3 650,000 1,900,000
4 600,000 2,500,000

cash outflow = 1,000,000

Pay back period = 1+ (1,000,000-300,000)/950,000

=1yr + 700,000/950,000= 1 yr+ 0.73 (or 73/12)

payback period = 1 year 6 months

Calculation of Payback period of F1-11

Year Cashflow Cumulative cash flow
1 200,000 200,000
2 600,000 800,000
3 1,200,000 2,000,000
4 1,400,000 3,400,000

cash outflow is 1,200,0000

Payback period = 2 year + (1,200,000 cash outflow-800,000 cash inflow )/1,200,000(3rd year cash flow)

= 2 years + 400,000/1,200,000

=2 year + 0.33 (or 33/12)

=2 years and 2 months

Yes, after calculation of payback period of both the projects Nimrod payback period is less than F1-11, so, it would wise to choose project F1-11

c) some managers prefer the payback period method to NPV because,Payback period method tells how long it will take the project to recover its Investment or Break even stage.

Another reason choosing Payback period over NPV is, the estimation of discount rate is unsure or not having any exact method to determine cost of capital(discounted rate)